T Twenty years from now, you want to spend $175,000 for a fancy car. How much...

T

Twenty years from now, you want to spend $175,000 for a fancy car. How much must you deposit as a lump sum today to achieve this goal at an annual interest rate of 6.6 percent?

$54,208.16

$48,740.95

$57,911.08

$40,019.82

$51,446.60
Answer #1

Answer:

Amount required after 20 years (FV) = $175,000
Rate (r ) = 6.60%
Time (n) = 20 years
Amount to be deposited today (PV) = ??

Future Value = Present Value * (1 + r)^ n
$175,000 = PV * (1 + 0.0660) ^ 20
$175,000 = PV * 1.0660^ 20
$175,000 = PV * 3.59041
PV = $48,740.95

Therefore, an amount of $48,740.95 is required to be deposited today to achieve the goal.